what is the equation of the line is slope-intercept form is y=___. (-1,-3) and (-8,-9)
Answer by mananth(16949)

x1 y1 x2 y2
-1 -3 -8 -9

slope m = (y2-y1)/(x2-x1)
( -9 - -3 )/( -8 - -1 )
( -6 / -7 )
m= 6/ 7

Plug value of the slope and point ( -1 , -3 ) in
Y = m x + b
-3.00 = - 6/7 + b
b= -3 - - 6/7
b= -15/ 7
So the equation will be
Y = 6/7 x -15/7
m.ananth@hotmail.ca